Curriculum Vitae

ADEBISI AGBOOLA

Department of Mathematics  Office: (805) 893-3844
University of California  Dept.: (805) 893-2171
Santa Barbara, CA 93106 Fax: (805) 893-2385
e-mail: agboola@math.ucsb.edu  Home: (805) 685-3050

url: http://www.math.ucsb.edu/˜ agboola

Research Interests:

Number Theory, Arithmetic Algebraic Geometry.

Regular Faculty Positions:

• University of California, Santa Barbara (July 2002 – present) Professor

• University of California, Santa Barbara (July 1998 – June 2002) Associate Professor

• University of California, Santa Barbara (July 1995 – June 1998) Assistant Professor

• University of California, Berkeley (July 1992 – June 1995) Charles B. Morrey Assistant Professor

Visiting Positions:

• CIRM, Université de Montreal, Montreal (August 2005 – December 2005) Visiting Researcher

• Harvard University, Cambridge (Sept. 1999 – August 2000) Visiting Scholar

• Université de Bordeaux I, Bordeaux (May 1998 – June 1998) Professeur Invité

• Institute Henri Poincaré, Paris (June – July 1997) Member

• Institute For Advanced Study, Princeton (Sept. 1995 – August 1996) Member

• Mathematical Sciences Research Institute, Berkeley (Sept. 1991 – August 1992)

Postdoctoral Research Fellow Education:

• Columbia University, New York Ph.D. in Mathematics (May 1991) Adviser: T. Chinburg

• Columbia University, New York M.A. in Mathematics (May 1988)

• University of Cambridge, Cambridge, England Certificate of Advanced Study in Mathematics (with Distinction) (June 1986)

• University of Cambridge, Cambridge, England B.A. (Hons.) in Mathematics (June 1985)

Grants, Honors and Awards:

• NSF Grant Principal Investigator, ‘Galois Structure, Iwasawa Theory and Arithmetic geometry’ (NSF DMS-040139) (2004–2007)

• NSF Grant Principal Investigator, ‘Arithmetic Geometry and Galois Module Theory’ (NSF DMS-0070449) (2000–2004)

• NSF Grant Principal Investigator, ‘Iwasawa Theory and Arithmetic Geometry’ (NSF DMS-9700937) (1997–2001)

• NSF Mathematical Sciences Postdoctoral Fellowship (1991–1994)

• Columbia University Faculty Fellowship (1987–91)

• College Examination Prize, University of Cambridge (1986)

• College Scholarship, University of Cambridge (1985–86)

Service:

• Reviewer of NSF grants.

• Referee for: Annals of Mathematics, Mathematische Annalen, Crelle’s Journal, Journal of Number Theory, London Mathematical Society, Contemporary Mathematics, AMS Transactions, Compositio Mathematica, Documenta Mathematica, Quarterly Journal of Pure and Applied Mathematics, Annales de L’Institut Fourier.

• Department Colloquium Chair, 2001–02.

• Co-organiser, Arithmetic and Geometry Seminar, 1997–present.

• Departmental Library Liason, 1997–98, 2001–03, 2003–04, 2006–present.

• Member, University Southern Regional Library Facility Committee, 1998–99.

• Member, Departmental Graduate Committee, 1997–98, 2000–01.

• Member, Departmental Undergraduate Committee, 1996–97.

• Member, Departmental Recruitment Committee, 1997–present.

• Member, Departmental CCS Committee, 2001–present.

• Chair, Departmental CCS Committee, Fall 2006.

• Member, CCS Executive Committee, Fall 2006.

• Member, Departmental Personnel Committee, 2006–07.

• Departmental MSRI Representative, 2004–present.

• Co-organiser, AMS Special Session on ‘Arithmetic Geometry’, UCSB, April 2005.

• Chair, Departmental CCS Committee, 2007–2008.

• Member, CCS Executive Committee, 2007–2008.

Personal Data:

Born August 11, 1964, in Ogbomoso, Nigeria.

U.S. Citizen.

Bibliography:

1. Abelian varieties and Galois module structure in global function fields, Columbia University Thesis, New York, 1991.

2. Iwasawa Theory of elliptic curves and Galois module structure, Duke Math. J., 71, (1993), 441–462.

3. Abelian varieties and Galois module structure in global function fields, Math. Zeit., 217, (1994), 407–419.

4. Class invariants of Mordell-Weil groups, (with M. J. Taylor), J. Reine Agnew Math., 447, (1994), 23–61.

5. A geometric description of the class invariant homomorphism, J. de Th. des Nombres de Bordeaux, 6, (1994), 273–280.

6. Torsion points on elliptic curves and Galois module structure, Invent. Math., 123, (1996), 105–122.

7. A note on elliptic curves and Galois module structure in global function fields, American J. Math., 118, (1996), 427–438.

8. On p-adic height pairings and locally free classgroups of Hopf orders, Math. Proc. Cam. Phil. Soc., 123, (1998), 447–459.

9. On the Galois structure of equivariant line bundles on curves, (with D. Burns), American J. Math., 120, (1998), 1121–1163.

10. Line bundles, rational points, and ideal classes, (with G. Pappas), Math. Res. Letters, 7, (2000), 709–717.

11. On primitive and realisable classes, Compositio Mathematica, 126, (2001), 113–122.

12. On arithmetic class invariants, (with G. Pappas), Math. Annalen, 320, (2001), 339–365.

13. Grothendieck groups of vector bundles on schemes over finite fields, (with D. Burns), K-theory, 23, (2001), 251–303.

14. Galois modules and p-adic representations, preprint.

15. Twisted forms and relative algebraic K-theory, (with D. Burns), Proc. London Math. Soc. 92 (2006), 1–28.

16. Anticyclotomic Iwasawa theory of CM elliptic curves, (with B. Howard, and with an appendix by K. Rubin), Annales, de L’Institut Fourier 56 (2006), 1001–1048.

17. Anticyclotomic Iwasawa theory of CM elliptic curves II, (with B. Howard), Math. Res. Letters 12 (2005), 611–621.

18. On Rubin’s variant of the p-adic Birch and Swinnerton-Dyer Conjecture, to appear in Compositio Mathematica.

Selected Seminars and Talks: 1990 & 1991 Université de Bordeaux I, Number theory seminar (September 1990) CIRM, Luminy, France, Conference on ‘Structures Galoisiennes arithmétiques’ (September 1990) Ohio State University, Workshop on ‘The arithmetic of function fields’ (June 1991) 1992 University of Minnesota, Number theory seminar (January 1992) Louisiana State University, Conference on ‘Quadratic forms and number theory’ (February 1992) University of California, Santa Cruz, Number theory seminar (March 1992) Sundance Institute, Utah, Conference of ‘Arithmetic algebraic geometry’ (May 1992) 1993 Irsee, Germany, Conference on ‘Integral Galois structures’ (July 1993) University of Minnesota, Colloquium and number theory seminar (October 1993) Arizona State University, Number theory seminar (December 1993) 1994 & 1995 Fields Institute, Waterloo, Canada, Workshop on ‘Galois module structure’ (February 1994) University of California, Berkeley/MSRI, Workshop on ‘Déssins d’enfants’ (April 1994) AMS special session, Arithmetic geometry, Chicago (March 1995) Mathematisches Forschungsinstitut Oberwolfach, Germany, Conference on ‘Orders in arithmetic’

(April 1995) University of Pennsylvania, Workshop on ‘Galois theory’ (October 1995) 1996 Institute for Advanced Study, Princeton, Number theory seminar (March 1996) University of Maryland, Number theory seminar (April 1996) Boston University, Number theory seminar (April 1996) Columbia University, Number theory seminar (May 1996) University of California, Irvine, Number theory seminar (November 1996)

1997 & 1998 California Institute of Technology, Number theory seminar (April 1997) AMS special session, Number theory and arithmetic geometry, Montreal (September 1997) University of Southern California, Number theory seminar (April 1998) Université de Bordeaux I, Several number theory seminars (May–June 1998) Anogia, Greece, Conference on ‘p-adic representations in arithmetic’ (July 1998) 1999 & 2000 Mathematisches Forschungsinstitut Oberwolfach, Germany, Conference on ‘Orders in arithmetic’

(January 1999) State University of New York, Albany, Colloquium (October 1999) Boston University, Number theory seminar (February 2000) McGill University, Number theory seminar (March 2000) Harvard University, Number theory seminar (March 2000) Columbia University, Number theory seminar (May 2000) Ohio State University, Number theory seminar (May 2000) University of Illinois, Urbana, Millenial conference on number theory (May 2000) 2001 & 2002 University of California, Irvine, Southern California algebraic geometry conference (February

2001) University of Rome I, Number theory seminar (March 2001) University of Lille, Conference on ‘Galois modules in arithmetic geometry’ (July 2001) Johns Hopkins University, Number theory seminar (September 2001) Mathematisches Forschungsinstitut Oberwolfach, Germany, Conference on ‘Orders in arithmetic’

(February 2002) California Institute of Technology, Number theory seminar (May 2002) Johns Hopkins University, Conference on ‘Stark’s conjectures and related topics’ (August 2002) University of Washington, Number theory seminar (November 2002)

2003 & 2004 University of Michigan, Number theory seminar (April 2003) University of California, Irvine, Number theory seminar (May 2003) Mathematisches Forschungsinstitut Oberwolfach, Germany, Conference on ‘Algebraic number the-

ory’ (July 2003) Banff International Research Station, Canada, Conference on ‘Current trends in arithmetic and

geometry’ (August 2003) University of Durham, Durham, England, NoMaDs seminar (September 2003) University of California, Berkeley, Number theory seminar (October 2003) University of California, Los Angeles, Number theory seminar (October 2003) University of Chicago, Number theory seminar (October 2003) Vanderbilt University, Seminar (October 2003) Stanford University, Number theory seminar (February 2004) Université de Bordeaux I, Number theory seminar (November 2004) 2005 & 2006 UC San Diego, Southern California Number Theory Day (April 2005) Boston University, Open questions and recent developments in Iwasawa theory, in honour of Ralph

Greenberg’s 60th birthday (June 2005) McGill University, Quebec-Vermont number theory seminar (December 2005) McMaster University, Number theory seminar (December 2005) University of Utah, AMS special session on ‘Arithmetic Geometry’ (October 2006) Caltech, Number theory seminar (December 2006) 2007 & 2008 University of Washington, Number theory seminar (March 2007) University of Arizona, AMS special session on ‘Arithmetic Geometry’ (April 2007)

Teaching:

• University of California, Berkeley Fall 1992 Introduction to abstract algebra Fall 1993 Introduction to complex analysis Spring 1994 Topics in number theory (graduate course) Fall 1994 Honors introduction to abstract algebra; Introduction to calculus Spring 1995 Introduction to complex analysis

• University of California, Santa Barbara Fall 1996 Abstract algebra I (graduate course) Winter 1997 Abstract algebra II (graduate course); Calculus 3A Spring 1997 Calculus 3B; Introduction to p-adic analysis (graduate course) Fall 1997 Algebraic number theory I (graduate course) Winter 1998 Algebraic number theory II (graduate course); Introduction to complex variable Spring 1998 Introduction to modular forms (graduate course) Fall 1998 Abstract algebra I (graduate course); Introduction to analysis I Winter 1999 Abstract algebra II (graduate course); Introduction to analysis II Spring 1999 Abstract algebra III (graduate course) Fall 2000 Abstract algebra I (graduate course) Winter 2001 Abstract algebra II (graduate course); Topics in algebra I Spring 2001 Abstract algebra II (graduate course); Topics in algebra II Fall 2001 Complex variable I (graduate course); Topics in algebra I Winter 2002 Transition to higher mathematics Spring 2002 Transition to higher mathematics; Calculus 3A Fall 2002 Topics in algebra I; Introduction to number theory I Winter 2003 Introduction to number theory II Spring 2003 Transition to higher mathematics; Introduction to number theory III Fall 2003 Abstract algebra I (graduate course); Topics in algebra I Winter 2004 Abstract algebra II (graduate course) Spring 2004 Abstract algebra III (graduate course)

Winter 2005 Introduction to number theory I Spring 2005 Transition to higher mathematics; Introduction to number theory II Winter 2006 Number theory I (graduate course); CCS Introduction to analysis Spring 2006 Number theory II (graduate course); 34B (Calculus for life and social sciences) Fall 2006 CCS Introduction to higher mathematics Winter 2007 Elliptic curves I (graduate course) Spring 2007 Elliptic curves II (graduate course); Introduction to number theory